OpenStudy (anonymous):

Solve for x: −2(x + 3) = −2(x + 1) − 4 2 3 All real numbers No solution

OpenStudy (loser66):

Any idea??

OpenStudy (loser66):

no idea?

OpenStudy (loser66):

\[\huge -2(x+3)=-2(x+1)-4\\\\\huge-2x-6=-2x-2-4\\\\\huge -2x+6 =-2x-6\\\\\huge +2x~both~sides~6=-6\\\\\huge MAKE~SENSE?~NO!!\\\\\huge~Therefore~no~solution\]